OpenStudy (anonymous):

Evaluate the following expression for n = -3 and p = 4 3(2n – p) + 5p – 2(n + 3p)

OpenStudy (anonymous):

3(-6-4)+20-2(-3+12) -30+20+6-24=-4-24=-28

jimthompson5910 (jim_thompson5910):

Plug in n = -3 and p = 4 and evaluate 3(2n – p) + 5p – 2(n + 3p) 3(2(-3) – 4) + 5(4) – 2(-3 + 3(4)) 3(-6 – 4) + 20 – 2(-3 + 12) 3(-10) + 20 – 2(9) -30 + 20 – 18 -28 So 3(2n – p) + 5p – 2(n + 3p) = -28 when n = -3 and p = 4